OpenStudy (anonymous):

Write an equation of a line in point slope form that has a slope of -3 and passes through the point (3, -4)

OpenStudy (anonymous):

@Conqueror @satellite73 @Somy @ganeshie8 @sleepyjess

OpenStudy (anonymous):

do you know the "point slope" formula?

OpenStudy (anonymous):

i do but i always find it using slope intercept form ......can you see if my equation is right @satellite73

OpenStudy (anonymous):

use the point slope formula first, let me know what you get

OpenStudy (anonymous):

i got y=-3x+5 @satellite73

OpenStudy (anonymous):

did you start with \[y+4=-3(x-3)\]

OpenStudy (anonymous):

yes i guess so since your answer is right

OpenStudy (anonymous):

no i did -4=(-3)(3)+b then solved for b then just plugged it in for y=mx+b @satellite73